# MiCRo: Mixture Modeling and Context-aware Routing for Personalized Preference Learning

### v1

@@ Abstract
Jingyan Shen2thanks: Equal contribution. Emails: {jiarui14, ry21, yifan50, ruip4, tozhang, hanzhao}@illinois.edu, js5544@columbia.edu, fl38@rice.edu, Jiarui Yao [^1] footnotemark: 1, Rui Yang [^1] footnotemark: 1, Yifan Sun1, Feng Luo3, Rui Pan1, Tong Zhang1Han Zhao1 1University of Illinois at Urbana-Champaign, 2Columbia University, 3Rice University
@@ 3.1 Problem Setup
Let $\mathcal{X}$ denote the space of prompts, and $\mathcal{Y}$ denote the space of responses. Denote $\Delta^{K}=\{\mathbf{x}\in\mathbb{R}^{K}\mid\sum_{i=1}^{K}x_{i}=1,x_{i}\geq 0,i=1,\ldots,K\}$ as the $(K-1)$-dimensional probability simplex. In standard preference learning, human preferences are modeled based on the classic BT model. Specifically, for a given prompt $x\in\mathcal{X}$ and an LLM $\pi$, two candidate responses $a_{w},a_{l}\in\mathcal{Y}$ are sampled independently from $\pi(\cdot\mid x)$. The probability that a human annotator prefers $a_{w}$ over $a_{l}$ is given by:
@@ 3.2 Irreducible error of single BT model
We defer the detailed proof of Theorem 3.2 to Appendix A. In the lower bound, $s_{k}^{*}\sharp\pi(a_{w},a_{l}\mid x)$ is the induced distribution of $s^{*}_{k}$ acting on the pair of responses $(a_{w},a_{l})\sim\pi$ from the LLM given the prompt $x$, and the variance operator $\mathrm{Var}_{z}$ is applied with respect to the $K$ groups. $H(x,\pi,\mathbb{P}(z|x))$ is the Shannon entropy of the joint distribution over preference data given by the prompt, subpopulations, as well as the LLM $\pi$ (c.f. Appendix A for its definition). At a colloquial level, the lower bound says that, the more diverse the ground-truth scores $s_{k}^{*}$ from each subpopulation (hence a larger variance), or the subpopulation distribution $\mathbb{P}(z\mid x)$ (hence a larger $\rho$ and entropy), then the larger the cross-entropy loss of using a single BT model.
@@ 4.1 Mixture Modeling for Diverse Preferences
Unlike static and unconditional mixture approach used in previous work Chakraborty et al. (2024a), our formulation from Equation  (1) introduces a dynamic, context-aware weighting mechanism for mixture models by conditioning the subpopulation weights $\mathbb{P}(z=k\mid x)$ on the given prompt $x$. We emphasize that this is a crucial design that allows for contextual specialization, where prompts automatically activate the most relevant subpopulation’s reward model. By mimicking real-world expertise allocation, our approach avoids the diluted performance of static averaging. We provide a more detailed comparison of our method with existing works in Table 1.
@@ 5.1 Experimental Setup
We train the mixture reward models on binary-labeled preference datasets: HelpSteer2 Wang et al. (2024c), RPR Pitis et al. (2024), and preference-700K Dong et al. (2024). HelpSteer2 and RPR datasets contain human-labeled response pairs evaluated across multiple assessment dimensions. We construct the binary-labeled sets with the following process. For each dimension, we extract binary preference pairs based on absolute ratings, treating responses with higher ratings as “chosen” and those with lower ratings as “rejected.” Pairs with identical ratings are excluded from both training and test sets. To ensure diversity in preferences, we exclude pairs where all attributes are unanimously agreed upon from the training set. Ultimately, all pairs from all dimensions are mixed, resulting in 23.5K samples and 5.8K training samples from HelpSteer2 and RPR, respectively. We also apply the mixture model training on preference-700K dataset, a large-scale pairwise dataset created by aggregating instructions and response pairs from multiple data sources. Further details on these datasets are provided in Appendix D.2.
@@
We use the best 3B open-source reward model GRM-Llama3.2-3B Yang et al. (2024a) as the backbone, keeping it frozen while training $K$ linear probing heads on top. The router network is implemented as a one-layer MLP containing 128 units and a softmax activation. We provide full implementation details in Appendix D.1.
Baselines We evaluate the following baselines: 1) Single Reward: A single-head model trained using the standard BT loss. 2) Static Mixture: A simplified variant of our method, corresponding to the approach used in MaxMin-RLHF Chakraborty et al. (2024a), where the mixture model is trained with fixed, input-independent weights, without leveraging contextual information. 3) Shared-Base Ensemble Model: Lee et al. (2024b) introduces HyRe, a test-time adaptation approach based on ensemble networks. We adopt the multi-head architecture with a frozen prior network and multiple trainable heads, optimizing a uniformly weighted sum of BT losses. 4) Fully Supervised Model: We include ARMO Wang et al. (2024b), an 8B model trained on more than 500K fine-grained labels, as a baseline with full supervision.
@@ 5.2 Stage-1 Evaluation: Can MiCRo Disentangle Diverse Human Preferences?
In this experiment, we train MiCRo and baseline models (except ARMO) on HelpSteer2 and RPR training sets. We then evaluate the learned heads on the HelpSteer2 and RPR test sets, which cover 14 distinct preference dimensions. Each head is evaluated individually on every dimension. To ensure fair comparisons, we use the same number of heads $K$ for MiCRo and other multi-head baselines. The full evaluation results of MiCRo mixture heads are provided in Appendix E.
@@
Training Dataset Method Clarity and Conciseness Creativity and Originality Scientific Rigor User- Friendliness Narrative and Storytelling Quality Pedagogical Effectiveness Linguistic Creativity Factual Accuracy Humor and Entertainment Average preference Single Head 0.8679 0.6806 0.9286 0.6067 0.6500 0.8065 0.8077 0.9155 0.9405 0.8004 Shared-base Best Head 0.8302 0.8056 0.8095 0.8089 0.6500 0.7903 0.8654 0.8169 0.9167 0.8009 -700K Static Mixture 0.8679 0.6250 0.9048 0.6292 0.6500 0.8065 0.7981 0.9014 0.9286 0.7902 MiCRo (Ours) 0.9358 0.6833 0.9190 0.6764 0.6700 0.7484 0.7827 0.9380 0.9405 0.8105 HelpSteer2 Single Head 0.8491 0.6667 0.9048 0.6180 0.6500 0.8065 0.7404 0.8732 0.8690 0.7753 Shared-base Best Head 0.8491 0.6667 0.9048 0.6629 0.7000 0.8065 0.7404 0.8310 0.9048 0.7851 Static Mixture 0.9057 0.6389 0.9048 0.6854 0.6250 0.7903 0.7404 0.8451 0.9167 0.7838 MiCRo (Ours) 0.9245 0.6667 0.8690 0.7079 0.6875 0.8226 0.7692 0.8592 0.9286 0.8133
@@ 5.3 Stage-2 Evaluation: Can MiCRo Adapt to Personalized Preference?
The RPR training dataset includes user-specific criteria for each preference pair, explicitly specifying the user’s intent and evaluation dimension, and thus provides a well-defined source of user context. For HelpSteer2, we follow the approach of Pitis et al. (2024) and augment generic prompts with attribute-specific modifications based on the original assessment dimensions in the annotation process Wang et al. (2024c). Examples of contexts are provided in Appendix D.2. For training and test datasets, we prepend the contextual information to the user prompt and fine-tune the router accordingly.
@@ Results
As shown in Tab.2 and Tab.3, MiCRo achieves average test accuracies of 0.7830 on HelpSteer2 and 0.8218 on RPR under the within-dataset evaluation setting, outperforming all three baselines trained with binary labels. This highlights the effectiveness of the router in adapting to specific user preferences. The relatively lower performance in certain attributes can be attributed to noisy or inconsistent contextual signals.
Compared to methods requiring stronger supervision, MiCRo performs competitively with ARMO on RPR and outperforms it on HelpSteer2. Furthermore, we find that applying test-time adaptation to MiCRo ’s mixture heads outperforms the original HyRe, indicating that our first-stage training provides a stronger base without requiring explicit supervision. While HyRe benefits from test-time labels, it assumes access to labeled examples for each user at inference time. Context-aware routing offers a more practical alternative by generalizing to unseen users, but its performance depends on the quality of the context. In practice, we believe a richer and more informative context could further boost its effectiveness. Our ablation in Section 5.4 further shows that performance improves with access to more context samples. In general, these results highlight MiCRo as a practical and label-efficient solution to learn personalized preferences.
Tab. 4 presents results under the unseen user setting, showing that MiCRo consistently outperforms other baselines trained with binary labels. This further demonstrates the router can generalize across user distributions with contextual information. Additional results on the RewardBench benchmark are provided in Appendix E.
@@ 5.4 Ablation Study
Fig. 4 shows the convergence of context-aware routing in Stage 2 on the RPR and HelpSteer2 test sets as the number of context-labeled samples per attribute increases. At budget 50 (i.e., 450 and 250 examples in total for each dataset, respectively), the average accuracy across 9 attributes on RPR test set increases sharply from around 0.705 to over 0.841, while the accuracy on HelpSteer2 plateaus around 0.785. In both cases, performance improves steadily with larger budgets, with most gains occurring early, demonstrating that the router can efficiently adapt using only a small number of contextual examples.
@@ 7 Limitations
Although our formulation is general, there is a limited availability of public datasets that provide rich and consistent user context information, making it difficult to comprehensively evaluate personalization capabilities. Our current implementation relies on access to explicitly defined context criteria and partially synthetic settings to simulate user-specific signals. However, in many real-world scenarios, user intent is often implicit—e.g., reflected in multi-turn dialogue, demographic metadata, or behavioral patterns. Incorporating such implicit user contexts into the routing process remains an important direction for future work.
@@ 8 Ethics Statement
The paper introduces a two-stage, context-aware mixture modeling framework that uses large-scale binary preference datasets to improve personalized preference learning. All experiments are conducted using publicly available models and datasets. The licenses for all datasets are listed in Appendix D.2, and we ensure complete compliance with all the license terms. While MiCRo improves scalability in personalized preference learning, the mixture heads are trained without explicit supervision. As a result, the learned mixture heads may encode preferences that are either beneficial or harmful, with a risk of inadvertently modeling undesirable or malicious intent. We therefore emphasize the importance of thorough evaluation and safety auditing to mitigate potential misuse.
@@ Acknowledgments
This work is supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies. Additionally, we thank Hanyang Chen for his assistance and helpful input.
@@
To ease the reading, we first restate Theorem 3.2 and provide its proof. See 3.2
@@
Appendix B Reproducibility
Our code repository is available at https://github.com/MaxwellJryao/MiCRo.
Appendix C Discussions
C.1 Choice of K
In this section, we empirically investigate how different choices of $K$ affect performance. As shown in Fig. 5, model performance remains stable as $K$ increases, suggesting that overestimating $K$ is relatively benign, redundant heads tend to receive low weights, and the best-performing head remains consistent. However, when $K$ is underestimated, the model suffers from misspecification, leading to degraded performance in the first stage. While a larger $K$ improves representation capacity, we observe that it can make convergence more difficult in the second-stage router training, as shown in Fig. 6. To mitigate this, a simple pruning strategy can be applied to remove heads with highly correlated predictions on a hold-out set, reducing redundancy without compromising accuracy.
(a) HelpSteer2
(b) RPR
Appendix D Experimental Details
D.1 Implementation Details
@@ D.2 Models and Datasets
The HelpSteer2 dataset contains human-labeled response pairs evaluated across five assessment dimensions: helpfulness, correctness, complexity, coherence, and verbosity. We include a summary of dataset statistics for HelpSteer2 and RPR datasets in Tab. 5(b).
@@
We listed an example of criterion in RPR dataset and the generated prompts for HelpSteer2 are listed in Tab. 6.
@@
Attribute Clarity Conciseness Creativity Originality Scientific Rigor User Friendliness Narrative Storytelling Pedagogical Effectiveness Linguistic Creativity Factual Accuracy Humor
@@
Appendix E Additional Experiment Results
In this section, we present additional experiment results.
Additional experimental results
We present a full evaluation of mixture heads on RPR datset and HelpSteer2 dataset in Tab. 7 and Tab. 8, which further demonstrate the diversity and benefits of mixture heads compared with the single reward. We also report results on the RewardBench benchmark in Tab. 9, showing improvements over the single-head baseline.
@@ Appendix E Additional Experiment Results
Model Helpfulness Correctness Coherence Complexity Verbosity Average
Single Reward 0.7838 0.6686 0.6914 0.7907 0.8816 0.7632
MiCRo Head 1 0.8108 0.7151 0.7407 0.7132 0.8289 -
MiCRo Head 2 0.8270 0.7035 0.7407 0.7209 0.8355 -
MiCRo Head 3 0.6595 0.6105 0.6543 0.8217 0.9276 -
MiCRo Head 4 0.6378 0.5523 0.5679 0.8217 0.9211 -
MiCRo Head 5 0.8108 0.7326 0.7469 0.7287 0.8487 -
MiCRo (Ours) 0.8324 0.7140 0.7543 0.7627 0.8513 0.7830
Table 9: Accuracy on RewardBench test set. We train the mixture heads on a combined dataset consisting of HelpSteer2 and the PKU-SafeRLHF dataset. MiCRo also outperforms the single reward model.
Model Chat Chat-hard Safety Reasoning Average
Single Reward (3B) 0.9693 0.6930 0.9135 0.9189 0.8737
MiCRo-Hedge (Ours) 0.9497 0.7544 0.9122 0.9322 0.8871

### v2

@@ Abstract
Jingyan Shen2thanks: Equal contribution. Emails: {jiarui14, ry21, yifan50, ruip4, tozhang, hanzhao}@illinois.edu, jingyan.s@nyu.edu, fl38@rice.edu. Code is available at https://github.com/uiuctml/MiCRo., Jiarui Yao [^1] footnotemark: 1, Rui Yang [^1] footnotemark: 1, Yifan Sun1, Feng Luo3, Rui Pan1, Tong Zhang1, Han Zhao1 1University of Illinois Urbana-Champaign, 2New York University, 3Rice University.
@@ 3.1 Problem Setup
Let $\mathcal{X}$ denote the space of prompts, and $\mathcal{Y}$ denote the space of responses. Denote $\Delta^{K}=\{\mathbf{x}\in\mathbb{R}^{K}\mid\sum_{i=1}^{K}x_{i}=1,x_{i}\geq 0,i=1,\ldots,K\}$ as the (K1)(K-1)-dimensional probability simplex. In standard preference learning, human preferences are modeled based on the classic BT model. Specifically, for a given prompt $x\in\mathcal{X}$ and an LLM $\pi$, two candidate responses $a_{w},a_{l}\in\mathcal{Y}$ are sampled independently from $\pi(\cdot\mid x)$. The probability that a human annotator prefers $a_{w}$ over $a_{l}$ is given by:
@@ 3.2 Irreducible error of single BT model
We defer the detailed proof of Theorem˜3.2 to Appendix A. In the lower bound, $s_{k}^{*}\sharp\pi(a_{w},a_{l}\mid x)$ is the induced distribution of $s^{*}_{k}$ acting on the pair of responses $(a_{w},a_{l})\sim\pi$ from the LLM given the prompt $x$, and the variance operator $\mathrm{Var}_{z}$ is applied with respect to the $K$ groups. $H(x,\pi,\mathbb{P}(z|x))$ is the Shannon entropy of the joint distribution over preference data given by the prompt, subpopulations, as well as the LLM $\pi$ (c.f. Appendix˜A for its definition). At a colloquial level, the lower bound says that, the more diverse the ground-truth scores $s_{k}^{*}$ from each subpopulation (hence a larger variance), or the subpopulation distribution $\mathbb{P}(z\mid x)$ (hence a larger $\rho$ and entropy), then the larger the cross-entropy loss of using a single BT model.
@@ 4.1 Mixture Modeling for Diverse Preferences
Unlike static and unconditional mixture approach used in previous work Chakraborty et al. (2024a), our formulation from Equation  (1) introduces a dynamic, context-aware weighting mechanism for mixture models by conditioning the subpopulation weights $\mathbb{P}(z=k\mid x)$ on the given prompt $x$. We emphasize that this is a crucial design that allows for contextual specialization, where prompts automatically activate the most relevant subpopulation’s reward model. By mimicking real-world expertise allocation, our approach avoids the diluted performance of static averaging. We provide a more detailed comparison of our method with existing works in Table˜1.
@@ 5.1 Experimental Setup
We train the mixture reward models on binary-labeled preference datasets: HelpSteer2 Wang et al. (2024c), RPR Pitis et al. (2024), and preference-700K Dong et al. (2024). HelpSteer2 and RPR datasets contain human-labeled response pairs evaluated across multiple assessment dimensions. We construct the binary-labeled sets with the following process. For each dimension, we extract binary preference pairs based on absolute ratings, treating responses with higher ratings as “chosen” and those with lower ratings as “rejected.” Pairs with identical ratings are excluded from both training and test sets. To ensure diversity in preferences, we exclude pairs where all attributes are unanimously agreed upon from the training set. Ultimately, all pairs from all dimensions are mixed, resulting in 23.5K samples and 5.8K training samples from HelpSteer2 and RPR, respectively. The preference-700K dataset is a large-scale pairwise dataset created by aggregating instructions and response pairs from multiple data sources. Further details on these datasets are provided in Appendix B.2.
@@
We use the best 3B open-source reward model GRM-Llama3.2-3B Yang et al. (2024a) as the backbone, keeping it frozen while training $K$ linear probing heads on top. The router network is implemented as a one-layer MLP containing 128 units and a softmax activation. Full implementation details are provided in Appendix B.1. We further evaluate with an open-source 8B reward model in Appendix C.5.
Baselines
We evaluate the following baselines: (1) Single Reward: A single-head model trained using the standard BT loss. (2) Static Mixture: A simplified variant of our method, corresponding to the approach used in MaxMin-RLHF Chakraborty et al. (2024a), where the mixture model is trained with fixed, input-independent weights, without leveraging contextual information. (3) Shared-Base Ensemble Model: Lee et al. (2024b) introduces HyRe, a test-time adaptation approach based on ensemble networks. We adopt the multi-head architecture with a frozen prior network and multiple trainable heads, optimizing a uniformly weighted sum of BT losses. (4) Fully Supervised Model: We include ARMO Wang et al. (2024b), an 8B model trained on more than 500K fine-grained labels, as a baseline with full supervision.
@@ 5.2 Stage-1 Evaluation: Can MiCRo Disentangle Diverse Human Preferences?
In this experiment, we train MiCRo and baseline models (except ARMO) on HelpSteer2 and RPR training sets. We then evaluate the learned heads on the HelpSteer2 and RPR test sets, which cover 14 distinct preference dimensions. Each head is evaluated individually on every dimension. To ensure fair comparisons, we use the same number of heads $K$ for MiCRo and other multi-head baselines. The full evaluation results of MiCRo mixture heads are provided in Appendix C.1.
Training Dataset Method Clarity Creativity Scientific Rigor User- Friendliness Storytelling Pedagogical Linguistic Creativity Factual Accuracy Humor Average preference Single Head 0.8679 0.6806 0.9286 0.6067 0.6500 0.8065 0.8077 0.9155 0.9405 0.8004 Shared-base Best Head 0.8302 0.8056 0.8095 0.8089 0.6500 0.7903 0.8654 0.8169 0.9167 0.8009 -700K Static Mixture 0.8679 0.6250 0.9048 0.6292 0.6500 0.8065 0.7981 0.9014 0.9286 0.7902 MiCRo (Ours) 0.9358 0.6833 0.9190 0.6764 0.6700 0.7484 0.7827 0.9380 0.9405 0.8105 HelpSteer2 Single Head 0.8491 0.6667 0.9048 0.6180 0.6500 0.8065 0.7404 0.8732 0.8690 0.7753 Shared-base Best Head 0.8491 0.6667 0.9048 0.6629 0.7000 0.8065 0.7404 0.8310 0.9048 0.7851 Static Mixture 0.9057 0.6389 0.9048 0.6854 0.6250 0.7903 0.7404 0.8451 0.9167 0.7838 MiCRo (Ours) 0.9245 0.6667 0.8690 0.7079 0.6875 0.8226 0.7692 0.8592 0.9286 0.8133
@@ 5.3 Stage-2 Evaluation: Can MiCRo Adapt to Personalized Preference?
The RPR training dataset includes user-specific criteria for each preference pair, explicitly specifying the user’s intent and evaluation dimension, and thus provides a well-defined source of user context. For HelpSteer2, we follow the approach of Pitis et al. (2024) and augment generic prompts with attribute-specific modifications based on the original assessment dimensions in the annotation process Wang et al. (2024c). Examples of contexts are provided in Appendix B.2. For training and test datasets, we prepend the contextual information to the user prompt and fine-tune the router accordingly.
@@ Results
As shown in Tab. 2 and Tab. 3, MiCRo achieves average test accuracies of 0.7830 on HelpSteer2 and 0.8218 on RPR under the within-dataset evaluation setting, outperforming all three baselines trained with binary labels. This highlights the effectiveness of the router in adapting to specific user preferences. The relatively lower performance in certain attributes can be attributed to the limited supervision budget, which may lead to a distribution mismatch between training and evaluation for some attributes. Our ablation in Appendix C.2 further shows that performance improves with access to more context samples. In practice, providing a richer and more informative context could further enhance the router’s performance.
Compared to methods requiring stronger supervision, MiCRo performs competitively with ARMO on RPR and outperforms it on HelpSteer2. Furthermore, we find that applying test-time adaptation to MiCRo ’s mixture heads outperforms the original HyRe, indicating that our first-stage training provides a stronger base without requiring explicit supervision. While HyRe benefits from test-time labels, it assumes access to labeled examples for each user at inference time. The context-aware routing offers a more practical alternative by generalizing to unseen users. In general, these results highlight MiCRo as a practical and label-efficient solution to learn personalized preferences.
Tab. 4 presents results under the unseen user setting, showing that MiCRo consistently outperforms other baselines trained with binary labels. This further demonstrates the router can generalize across user distributions with contextual information. Additional results on the RewardBench benchmark are provided in Appendix C.
@@ 5.4 Ablation Study
Fig. 4 shows the convergence of context-aware routing in Stage 2 on the RPR and HelpSteer2 test sets as the number of context-labeled samples per attribute increases. At budget 50 (i.e., 450 and 250 examples in total for each dataset, respectively), the average accuracy across 9 attributes on RPR test set increases sharply from around 0.705 to over 0.841, while the accuracy on HelpSteer2 plateaus around 0.785. In both cases, performance improves steadily with larger budgets, with most gains occurring early, demonstrating that the router can efficiently adapt using only a small number of contextual examples. More case studies on specific attributes can be found in Appendix C.2.
@@ 7 Limitations
Although our formulation is general, there is a limited availability of public datasets that provide rich and consistent user context information, making it difficult to comprehensively evaluate personalization capabilities. Our current implementation relies on access to explicitly defined context criteria and partially synthetic settings to simulate user-specific signals. However, in many real-world scenarios, user intent is often implicit, e.g., reflected in multi-turn dialogue, demographic metadata, or behavioral patterns. Incorporating such implicit user contexts into the routing process remains an imperative direction for future work.
@@ 8 Ethics Statement
The paper introduces a two-stage, context-aware mixture modeling framework that uses large-scale binary preference datasets to improve personalized preference learning. All experiments are conducted using publicly available models and datasets. The licenses for all datasets are listed in Appendix B.2, and we ensure complete compliance with all the license terms. While MiCRo improves scalability in personalized preference learning, the mixture heads are trained without explicit supervision. As a result, the learned mixture heads may encode preferences that are either beneficial or harmful, with a risk of inadvertently modeling undesirable or malicious intent. We therefore emphasize the importance of thorough evaluation and safety auditing to mitigate potential misuse. In light of this, we outline two safeguards to ensure safer deployment: (1) safety auditing with existing benchmarks: the mixture heads can be audited using safety-focused datasets such as PKU-SafeRLHF Ji et al. (2024a) and Anthropic HH Bai et al. (2022). If a head assigns unusually high scores to rejected responses in the benchmarks, the head should be masked during downstream use; (2) human-in-the-loop evaluation: human annotators can be involved to qualitatively inspect the high-reward outputs of individual heads to assess whether they reflect manipulative or unsafe patterns. With MiCRo ’s staged training setup, these pruning and masking steps do not interfere with Stage 2 router learning, allowing us to maintain performance while mitigating safety risks.
@@ Acknowledgments
This work is supported by NSF IIS grant No. 2416897 and No. 2442290, and an ONR grant No. N000142512318. HZ would like to thank Google for the support of a Google Research Scholar Award. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies. Additionally, we thank Hanyang Chen for his assistance and helpful input.
@@
To ease the reading, we first restate Theorem˜3.2 and provide its proof. See 3.2
@@
Appendix B Experimental Details
B.1 Implementation Details
@@ B.2 Models and Datasets
The HelpSteer2 dataset contains human-labeled response pairs evaluated across five assessment dimensions: helpfulness, correctness, complexity, coherence, and verbosity. We include a summary of dataset statistics for HelpSteer2 and RPR datasets in Table 5(b).
@@
We listed an example of criterion in RPR dataset and the generated prompts for HelpSteer2 are listed in Table 6.
@@
Attribute
Clarity
Conciseness
Creativity
Originality
Scientific
Rigor
User
Friendliness
Narrative
Storytelling
Pedagogical
Effectiveness
Linguistic
Creativity
Factual
Accuracy
Humor
@@
Appendix C Additional Experimental Results
C.1 Full evaluations of mixture heads
We present a full evaluation of mixture heads on RPR datset and HelpSteer2 dataset in Table 7 and Table 8, which further demonstrate the diversity and benefits of mixture heads compared with the single reward.
@@ C.1 Full evaluations of mixture heads
Model Helpfulness Correctness Coherence Complexity Verbosity Average Single Reward 0.7838 0.6686 0.6914 0.7907 0.8816 0.7632 MiCRo Head 1 0.8108 0.7151 0.7407 0.7132 0.8289 - MiCRo Head 2 0.8270 0.7035 0.7407 0.7209 0.8355 - MiCRo Head 3 0.6595 0.6105 0.6543 0.8217 0.9276 - MiCRo Head 4 0.6378 0.5523 0.5679 0.8217 0.9211 - MiCRo Head 5 0.8108 0.7326 0.7469 0.7287 0.8487 - MiCRo (Ours) 0.8324 0.7140 0.7543 0.7627 0.8513 0.7830
C.2 Ablation study on sample budget
To support fast and lightweight deployment, we train the router using only 50 labeled preference pairs per attribute. While this setup is effective for most cases, limited supervision can lead to train–test mismatch on certain fine-grained attributes. In Tab. 9, we present a case study on three representative attributes from the RPR test set to analyze the impact of increasing the router’s training budget. We observe consistent improvements in both accuracy and stability as the sample budget increases. For example, on Linguistic Creativity, MiCRo surpasses the single-reward baseline when trained with 150 samples.
[…]
C.3 Ablation study on choice of K
In this section, we empirically investigate how different choices of $K$ affect performance. As shown in Fig. 5, model performance remains stable as $K$ increases, suggesting that overestimating $K$ is relatively benign, redundant heads tend to receive low weights, and the best-performing head remains consistent. However, when $K$ is underestimated, the model suffers from misspecification, leading to degraded performance in the first stage. While a larger $K$ improves representation capacity, we observe that it can make convergence more difficult in the second-stage router training, as discussed in Section 4.2. To mitigate this, a simple method is to merge heads with highly correlated predictions on a hold-out set, which effectively reduces model size without compromising accuracy.
(a) HelpSteer2
(b) RPR
C.4 Evaluation results on RewardBench benchmark
Tab. 10 reports results on the RewardBench benchmark, demonstrating improvements over the single-head baseline.
Model Chat Chat-hard Safety Reasoning Average Single Reward (3B) 0.9693 0.6930 0.9135 0.9189 0.8737 MiCRo-Hedge (Ours) 0.9497 0.7544 0.9122 0.9322 0.8871
Table 10: Accuracy on RewardBench test set. We train the mixture heads on a combined dataset consisting of HelpSteer2 and the PKU-SafeRLHF dataset. MiCRo consistently outperforms the single reward model.
C.5 Evaluation results on 8B reward model
To assess scalability, we further evaluate with an open-source 8B reward model (GRM-Llama3-8B) Yang et al. (2024a) on HelpSteer and RPR test sets. As shown in Tab. 11 and Tab. 12, MiCRo consistently outperforms the baselines in terms of average accuracy, demonstrating its robustness across model variants.
Method Clarity Creativity Scientific Rigor User- Friendliness Storytelling Pedagogical Linguistic Creativity Factual Accuracy Humor Average Single Head 0.5094 0.5972 0.3333 0.7865 0.8125 0.5484 0.7596 0.3380 0.9167 0.6224 Static Mixture 0.3962 0.5641 0.3214 0.7753 0.7375 0.6129 0.8462 0.3380 0.8571 0.6055 ARMO 0.9057 0.6806 0.9405 0.6966 0.7875 0.7903 0.9135 0.9014 0.9463 0.8403 MiCRo-Stage-1 0.9434 0.8750 0.9286 0.8202 0.8750 0.8387 0.9231 0.9577 0.9405 0.9002 MiCRo (8B, B=50) 0.8189 $\pm$ 0.0350 0.7750 $\pm$ 0.0283 0.8500 $\pm$ 0.0381 0.8225 $\pm$ 0.0260 0.8725 $\pm$ 0.0366 0.8065 $\pm$ 0.0177 0.8654 $\pm$ 0.0136 0.8620 $\pm$ 0.0372 0.9214 $\pm$ 0.0143 0.8438 $\pm$ 0.0077
Table 11: Accuracy on the RPR test set. MiCRo-Stage-1 denotes the accuracy achieved by the best-performing heads from Stage 1 mixture learning. For MiCRo, we report the “mean $\pm$ standard deviation” across 5 independent runs using randomly sampled $B$ training samples per attribute.
Method Helpfulness Correctness Coherence Complexity Verbosity Average Single Head 0.7636 0.7318 0.6909 0.7682 0.7818 0.7473 Static Mixture 0.7818 0.7364 0.7136 0.7455 0.7636 0.7482 ARMO 0.6919 0.6395 0.7593 0.7132 0.7500 0.7108 MiCRo-Stage-1 0.7864 0.7273 0.7318 0.8136 0.8364 0.7791 MiCRo (8B, B=50) 0.7864 $\pm$ 0.0000 0.7227 $\pm$ 0.0043 0.7242 $\pm$ 0.0021 0.7727 $\pm$ 0.0021 0.7712 $\pm$ 0.0119 0.7555 $\pm$ 0.0027
Table 12: Accuracy on the HelpSteer test set. MiCRo-Stage-1 denotes the accuracy achieved by the best-performing heads selected from Stage 1 mixture learning. For MiCRo, we report the “mean $\pm$ standard deviation” across 5 independent runs using randomly sampled $B$ training samples per attribute.